OpenStudy (anonymous):

graph the quadratic equation. label the vertex axis of symmetry and x-intercepts indentify the domain and range ofd the function y=-x^2+9

OpenStudy (anonymous):

Domain is \[(-\infty,-\infty)\]  Range is \[(-\infty,9]\] Vertex axis of symmetry is Y axis. X intercepts are -3 & 3.